OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase? 102π in3 106π in3 100π in3 108π in3

OpenStudy (anonymous):

easy but lengthy!!

OpenStudy (anonymous):

Volume of marbles = 6 ( 4/3 pi 1^3) = 8 pi Volume of vase up to the water = pi . 3^2 . 12 = 108 pi Volume of water alone = 108 pi - 8 pi = 100 pi

OpenStudy (anonymous):

^^ this is violation of the code of conduct ,,, we have to explain not to solve ..

OpenStudy (anonymous):

lol u kidding hashirr

OpenStudy (anonymous):

no i m serious .... it hinders the learning process, you must explain your steps :)

OpenStudy (anonymous):

ok maybe il also have to learn... thanks for the warning :/ but arent my steps above self explanatory... ? spoon feeding others is boring to me. I think you show the steps, and they will understand. Not this simple always but.... lol. thanks anyways :)

OpenStudy (anonymous):

^^ well i used to do that too ..... they warned me about that :) ...... no problem ... its my pleasure :)